OpenStudy (anonymous):

Evaluate the expression below when x = -4. 4x2 - 2x × 15

OpenStudy (anonymous):

is this question correct..??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:A. -56 B. 184 C. 376 D. 1,080

OpenStudy (anonymous):

4(-4)^2 - 2(-4) x 15 = 4(16)+8x15 =64+120 =184